Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by thephoenix » Sat Jan 23, 2010 1:03 am
If 2 different representatives are to be selected at random from a group of 10 employees
and if p is the probability that both representatives selected will be women, is p >0.5
(1) More than 50% of the 10 employees are women.
(2) The probability that both representatives selected will be men is less than 10%
.
Join the discussion
Source: — Problem Solving |

by trindadesn » Sat Jan 23, 2010 2:07 am
Is the answer E ?
Join the discussion

by thephoenix » Sat Jan 23, 2010 2:16 am
trindadesn wrote:Is the answer E ?
yes
pls show your calculation
Join the discussion

by trindadesn » Sat Jan 23, 2010 2:52 am
Ststement 1. More than 50% of the 10 employees are women.
Possible options for the 10 members are
M W
4 6
3 7
2 8
1 9

for 4 men and 6 women prob that both are women is 6/10 * 5/9 = 5/30. this is <0.5
for 2 men and 8 women prob …… is 8/10 * 7/9 = 28/45 this is >0.5

St.1 insuff.

Statement 2. The probability that both representatives selected will be men is less than 10%
take same options ( you will have to try with men greater than women too)

for 3 men and 7 women prob that both are men is 3/10 * 2/9 = 1/15 this is < 10%
and prob that both are women is 7/10 * 6/9 = 7/15 , this is < 0.5
for 2 men and 8 women prob that both are men is 2/10 * 1/9 = 1/45 this is < 10%
but prob that both are women is 8/10 * 7/9 = 28/45 this is >0.5

so St 2 also insufficient.

Combining both cases will result in same set of sols as case 2 as we have only taken options where the women are greater in no than men.

I hope you understand my explanations.
Join the discussion

by thephoenix » Sat Jan 23, 2010 6:23 pm
trindadesn wrote:Ststement 1. More than 50% of the 10 employees are women.
Possible options for the 10 members are
M W
4 6
3 7
2 8
1 9

for 4 men and 6 women prob that both are women is 6/10 * 5/9 = 5/30. this is <0.5
for 2 men and 8 women prob …… is 8/10 * 7/9 = 28/45 this is >0.5

St.1 insuff.

Statement 2. The probability that both representatives selected will be men is less than 10%
take same options ( you will have to try with men greater than women too)

for 3 men and 7 women prob that both are men is 3/10 * 2/9 = 1/15 this is < 10%
and prob that both are women is 7/10 * 6/9 = 7/15 , this is < 0.5
for 2 men and 8 women prob that both are men is 2/10 * 1/9 = 1/45 this is < 10%
but prob that both are women is 8/10 * 7/9 = 28/45 this is >0.5

so St 2 also insufficient.

Combining both cases will result in same set of sols as case 2 as we have only taken options where the women are greater in no than men.

I hope you understand my explanations.
thsnks !!!!
is this the only way to solve ??
Join the discussion